OpenStudy (anonymous):

A fossil ammonite has the form of a logarithmic spiral. Yolanda, now a paleontologist, measures the distances of loops #2 and #4 along a radial line from the center of the spiral. She finds that the distances are 6 centimeters and 13.5 centimeters. The fossil is in poor condition, and she cannot measure the distance of loop #3 from the center. But she can calculate it. What is the distance in centimeters? Do not include units in your answer.

OpenStudy (anonymous):

i know the equation is a(n)=a(1)+(n-1)*d a(3)=6+(3-1)*7.5 but that gives me 21, which cant b right

OpenStudy (anonymous):

i got 20

OpenStudy (anonymous):

but that stilt cant b right cuz how can the 3rd loop b bigger thanthe fourth loop?

OpenStudy (anonymous):

what do you get whn you set up your formula to = your 4th loop?

OpenStudy (anonymous):

using the equation a(n)=a(n-1)+d a(4)=6+7.5 it gives 13.5 which is the measurement for the 4th loop

OpenStudy (anonymous):

a(3)=(6-1)+7.5?????

OpenStudy (anonymous):

it was 9. but how did they get 9 is wat i want to know

OpenStudy (anonymous):

haha idk